# Neural Risk Limiting Dispatch in Power Networks: Formulation and Generalization Guarantees

### v1

@@ IV-C1 Neuron number
IV-C2 Sample size
Fig. 7 compares the normalized suboptimality of all methods with different sample sizes. Similar to the results in Fig. 6, the proposed method always outperforms the rest ones by achieving the lowest normalized suboptimality. Moreover, with an increasing sample size, the suboptimality of the proposed method decreases. According to our proposed generalization performance guarantee, i.e., Theorem 1, the PAC bound of our method gradually approaches zero with the increase of the sample size. Other methods do not display a similar trend in suboptimality.
@@ V Conclusions
This paper presents a learning-based approach to solve the network RLD problem with provably near-optimal performance. It begins by designing a data-driven formulation for the network RLD problem, which aims to learn a decision rule from historical data. This decision rule directly maps day-ahead auxiliary information to cost-effective dispatch decisions for the future delivery interval, so it can bypass the inherent “mismatch” issue of the widely adopted predict-then-optimize paradigm. We then propose neural RLD as a novel solution method for this data-driven formulation. It employs a L2-regularized neural network to learn the decision rule, thereby transforming the above data-driven formulation as a neural network training task with a specialized training loss. The gradient of this loss is explicitly derived, allowing for efficient completion of the training task using SGD. A generalization performance guarantee is further developed based on uniform convergence, which provides a PAC upper bound for the suboptimality of the neural RLD. Our numerical experiments on IEEE 5-bus, 118-bus, 300-bus, and 1354-bus test systems demonstrate the neural RLD’s desirable performance in terms of convergence, suboptimality, and computational efficiency. Simulations also confirm the asymptotic optimality of our method, implying that its suboptimality diminishes to zero with high probability as more samples are utilized.
The proposed neural RLD offers an effective solution for the two-stage network RLD problem. However, its applicability to the original multi-stage stochastic control formulation remains limited. Additionally, our theoretical guarantees are currently confined to neural networks composed solely of fully connected layers. In view of these limitations, our future work aims to extend the neural RLD to the multi-stage formulation and to establish theoretical guarantees for more sophisticated learning models.
@@
Proof of Proposition 1: The optimal objective of the dual problem of (II-A) is expressed as:
@@
Since (II-A) can be reformulated as a linear program, the strong duality holds, so
@@
Proof of Theorem 1: This proof is composed of the following steps: 1) prove the Lipschitz continuity of the RLD loss (6); 2) obtain an upper bound of the Rademacher complexity, and 3) derive the PAC upper bound given in Theorem 1.
@@ B-1 Lipschitz continuity of the RLD loss
Suppose $(\mathbf{g}^{\star})^{(1)}$ and $(\mathbf{g}^{\star})^{(2)}$ are the optimal solutions of (II-A) when $\mathbf{u}=\mathbf{u}^{(1)}$ and $\mathbf{u}=\mathbf{u}^{(2)}$, respectively. According to [Theorem 2.2] in [33], the following inequality holds because (II-A) can be transformed into a linear problem by introducing auxiliary variables:
@@
where $C_{\mathbf{g}}$ is a finite constant. Then, we have:
@@
Note (19) also gives the definition of $C_{\ell}$ in Theorem 1. When the L2-regularized neural network $\mathbf{h}$ is used to predict the dispatch decision $\mathbf{u}$, we have:
@@
The above inequality shows that the mapping $\psi(\cdot):\mathbf{h}(\mathbf{x})\mapsto\ell(\mathbf{h}(\mathbf{x}),\mathbf{d})$ is $C_{\ell}$-Lipschitz continuous.
@@ B-2 Upper bound of Rademacher complexity
The empirical Rademancher complexity of the loss class on dataset $\mathcal{S}$ is defined as follows:
^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})=\mathbb{E}_{\xi}\left[\sup_{\mathbf{\ell}\in\mathcal{L}_{\mathbf{h}}}\sum_{m\in\mathcal{M}}\xi^{(m)}\ell(\mathbf{h}(\mathbf{x}^{(m)}),\mathbf{d}^{(m)})\right].$ (23)
where ^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$ is calculated based on observed realizations, which differs from the Rademancher complexity $\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$. To obtain the upper bound of ^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$, we first introduce the vector contraction inequality in the following lemma:
Lemma 1 (Vector contraction inequality [33]).
@@
Then, this empirical Rademancher complexity is bounded by:
@@
where $\mathcal{X}=\{\mathbf{x}_{i},\forall m\in\mathcal{M}\}$. Observing that the empirical Rademancher complexity of the hypothesis class $\mathcal{H}_{j}$ on dataset $\mathcal{X}$ is defined as:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})=\mathbb{E}_{\xi}\left[\sup_{{h}_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\xi^{(m,j)}{h}_{j}(\mathbf{x}^{(m)})\right],$ (28)
@@
We further define a new variable $\overline{\mathbf{w}}^{k}_{j}=\frac{\mathbf{w}^{k}_{j}}{{\|\mathbf{w}^{k}_{j}\|_{2}}}$. Obviously, we have $\|\overline{\mathbf{w}}^{k}_{j}\|_{2}=1$. Then, according to the definition of $h_{j}(\mathbf{x})$, the empirical Rademancher complexity ^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$ can be expressed as:
@@
$\leq\frac{(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}=1,\end{subarray}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|\right]$ (36)
$\leq\frac{(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|\right],$ (39)
@@
$\leq\frac{2(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right]$ (42)
$\leq\frac{2(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\sum_{m\in\mathcal{M}}\xi^{(m)}\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right]$ (45)
@@
where the definition of class ¯$\overline{\mathcal{H}}^{k}$ is almost the same as ${\mathcal{H}}^{k}$ defined in (22) except for replacing the L2 regulation constraint $\|\mathbf{w}^{(k)}_{j}\|_{2}\leq W^{\mathrm{max}}$ with $\left\|\overline{\mathbf{w}}\right\|_{2}\leq 1$. The second inequality holds because $\phi(\cdot)$ is 1-Lipschitz continuous. The last equality holds due to the definition of the empirical Rademancher complexity. In a similar way, we can also obtain:
@@
The Rademancher complexity of the loss is the expectation of the empirical one over the probability distribution of $\mathbf{x}$. By substituting (48) into (29), we can bound this Rademancher complexity via:
@@
$\leq\sqrt{2}C_{\ell}\cdot\sqrt{W^{\mathrm{max}}}\cdot(2W^{\mathrm{max}})^{K-1}\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1}),$ (49)
where ¯$\overline{\mathcal{H}}^{1}=\{\mathbf{x}\rightarrow\overline{\mathbf{w}}^{\intercal}\mathbf{x},\|\overline{\mathbf{w}}\|_{2}\leq 1\}$. According to [Theorem 3] in [35], its Rademancher complexity is bounded by:
$\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1})\leq\frac{X^{\mathrm{max}}}{\sqrt{M}}.$ (50)
By substituting (50) into (49), we obtain an upper bound of the Rademancher complexity $\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$, as follows:
@@ B-3 PAC upper bound of the excess cost
Lemma 2 (Generalization bound via uniform convergence [36]).

### v2

@@ IV-C1 Neuron number
IV-C2 Hidden layer number
We further investigate the impact of the hidden layer number. While keeping the total neuron number fixed at 60, we vary the hidden layer number from 1 to 6, setting the neurons per layer to 60, 30, 20, 15, 12, and 10, respectively. The suboptimality results for the different methods are shown in Fig. 8. For the proposed method, a slight growing trend in suboptimality is observed as the hidden layer number increases. This trend results from the fact that, in our setup, shallower networks have more learnable parameters than deeper ones. [^7] The total number of learnable parameters for networks with 1, 2, 3, 4, 5, and 6 hidden layers are 14160, 7980, 5520, 4215, 3408, and 2860, respectively. If the NNs are well-trained, having more learnable parameters (as in the shallower networks) can reduce the generalization error. Thus, shallower networks may achieve lower suboptimality. Moreover, the proposed method directly employs the RLD objective as its training loss, inherently minimizing suboptimality. In contrast, benchmarks Imitation and Two-step use MSE as their training loss, which can not accurately measure suboptimality. Therefore, even when the MSE is minimized, these benchmarks may still exhibit higher suboptimality compared to the proposed method.
Figure 8: Normalized suboptimality of all methods with different hidden layer numbers in the 118-bus test case.
IV-C3 Sample size
Fig. 9 compares the normalized suboptimality of all methods with different sample sizes. Similar to the results in Fig. 7, the proposed method always outperforms the rest ones by achieving the lowest normalized suboptimality. Moreover, with an increasing sample size, the suboptimality of the proposed method decreases. According to our proposed generalization performance guarantee, i.e., Theorem 1, the PAC bound of our method gradually approaches zero with the increase of the sample size. Other methods do not display a similar trend in suboptimality. Moreover, our proposed method trains the NN by minimizing the RLD loss (6) without requiring labelled data. In other words, the NN is trained in a self-supervised manner. Thus, our approach enables the easy generation of synthetic day-ahead observable information to further enhance the performance of our method.
@@ V Conclusions
This paper presents a learning-based approach to solve the network RLD problem with provably near-optimal performance. It begins by designing a data-driven formulation for the network RLD problem, which aims to learn a decision rule from historical data. This decision rule directly maps day-ahead auxiliary information to cost-effective dispatch decisions for the future delivery interval, so it can bypass the inherent “mismatch” issue of the widely adopted predict-then-optimize paradigm. We then propose neural RLD as a novel solution method for this data-driven formulation. It employs a L2-regularized NN to learn the decision rule, thereby transforming the above data-driven formulation as a NN training task with a specialized training loss. The gradient of this loss is explicitly derived, allowing for efficient completion of the training task using SGD. A generalization performance guarantee is further developed based on uniform convergence, which provides a PAC upper bound for the suboptimality of the neural RLD. Our numerical experiments on IEEE 5-bus, 118-bus, 300-bus, and 1354-bus test systems demonstrate the neural RLD’s desirable performance in terms of convergence, suboptimality, and computational efficiency. Simulations also confirm the asymptotic optimality of our method, implying that its suboptimality diminishes to zero with high probability as more samples are utilized.
The proposed neural RLD provides an effective solution to the two-stage network RLD problem. However, its applicability to the original multi-stage stochastic control formulation remains limited. Furthermore, our current theoretical guarantees are restricted to NNs consisting solely of fully connected layers. To address these limitations, future work will focus on extending the neural RLD framework to the multi-stage formulation and developing theoretical guarantees for more advanced learning models.
@@
Proof of Proposition 1: This proposition is an application of the envelope theorem [38]. A similar idea has also been reported in [32]. Specifically, the optimal objective of the dual problem of (2) is expressed as:
@@
Since (2) can be reformulated as a linear program, the strong duality holds, so
@@
Proof of Theorem 1: Fig. 10 outlines the detailed steps for proving Theorem 1, as follows: i) Prove the Lipschitz continuity of the RLD loss (6) and calculate the corresponding Lipschitz constant, ii) Estimate the empirical and true Rademacher complexities of the loss class, and iii) Derive the PAC upper bound as stated in Theorem 1.
Figure 10: Detailed steps for proving Theorem 1, where RC represents Rademacher complexity.
@@ B-1 Lipschitz continuity of the RLD loss
The solution function of problem (2) maps the first-stage decision $\mathbf{u}$ and net demand $\mathbf{d}$ to the second-stage optimal solution $\mathbf{g}^{\star}$. According to [39, Theorem 2.2], this solution function is Lipschitz continuous, and the following inequality holds:
@@
where $\mathbf{u}^{(1)}$ and $\mathbf{u}^{(2)}$ are two different first-stage decisions, and $\mathbf{g}^{\star})^{(1)}$ and $(\mathbf{g}^{\star})^{(2)}$ are the corresponding optimal second-stage decisions. Parameter $C_{\mathbf{g}}$ is the Lipschitz constant of this solution function, and its detailed formulation is provided in [39, Theorem 2.2]. Then, we have:
@@
Note (19) also gives the definition of $C_{\ell}$ in Theorem 1. When the L2-regularized NN $\mathbf{h}$ is used to predict the dispatch decision $\mathbf{u}$, we have:
@@
The above inequality shows that the mapping $\psi(\cdot):\mathbf{h}(\mathbf{x})\mapsto\ell(\mathbf{h}(\mathbf{x}),\mathbf{d})$ is $C_{\ell}$-Lipschitz continuous. Then, according to
@@ B-2 Upper bound of Rademacher complexity
The empirical Rademacher complexity of the loss class on dataset $\mathcal{S}$ is defined as follows:
^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})=\mathbb{E}_{\xi}\left[\sup_{\mathbf{\ell}\in\mathcal{L}_{\mathbf{h}}}\sum_{m\in\mathcal{M}}\xi^{(m)}\ell(\mathbf{h}(\mathbf{x}^{(m)}),\mathbf{d}^{(m)})\right],$ (23)
where ^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$ is calculated based on observed realizations, which differs from the Rademacher complexity $\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$. To obtain the upper bound of ^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$, we first introduce the vector contraction inequality in the following lemma:
Lemma 1 (Vector contraction inequality [39])
@@
Then, this empirical Rademacher complexity is bounded by:
@@
where $\mathcal{X}=\{\mathbf{x}_{i},\forall m\in\mathcal{M}\}$. Observing that the empirical Rademacher complexity of the hypothesis class $\mathcal{H}_{j}$ on dataset $\mathcal{X}$ is defined as:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})=\mathbb{E}_{\xi}\left[\sup_{{h}_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\xi^{(m,j)}{h}_{j}(\mathbf{x}^{(m)})\right].$ (28)
@@
We further define a new variable $\overline{\mathbf{w}}^{k}_{j}=\frac{\mathbf{w}^{k}_{j}}{{\|\mathbf{w}^{k}_{j}\|_{2}}}$. Obviously, we have $\|\overline{\mathbf{w}}^{k}_{j}\|_{2}=1$. Then, according to the definition of $h_{j}(\mathbf{x})$, the empirical Rademacher complexity ^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$ can be expressed as:
@@
$\leq\frac{(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}=1,\end{subarray}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|\right]$
$\leq\frac{(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|\right].$ (34)
@@
$\leq\frac{2(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right]$
$\leq\frac{2(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\sum_{m\in\mathcal{M}}\xi^{(m)}\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right]$
@@
where the definition of class ¯$\overline{\mathcal{H}}^{k}$ is almost the same as ${\mathcal{H}}^{k}$ defined in (22) except for replacing the L2 regulation constraint $\|\mathbf{w}^{(k)}_{j}\|_{2}\leq W^{\mathrm{max}}$ with $\left\|\overline{\mathbf{w}}\right\|_{2}\leq 1$. The second inequality holds because $\phi(\cdot)$ is 1-Lipschitz continuous. The last equality holds due to the definition of the empirical Rademacher complexity. In a similar way, we can also obtain:
@@
The Rademacher complexity of the loss is the expectation of the empirical one over the probability distribution of $\mathbf{x}$. By substituting (37) into (29), we can bound this Rademacher complexity via:
@@
$\leq\sqrt{2}C_{\ell}\cdot\sqrt{W^{\mathrm{max}}}\cdot(2W^{\mathrm{max}})^{K-1}\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1,L2}),$ (38)
where ¯$\overline{\mathcal{H}}^{1,L2}=\{\mathbf{x}\rightarrow\overline{\mathbf{w}}^{\intercal}\mathbf{x},\|\overline{\mathbf{w}}\|_{2}\leq 1\}$. According to [41, Theorem 3], its Rademacher complexity is bounded by:
$\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1,L2})\leq\frac{X^{\mathrm{max}}}{\sqrt{M}}.$ (39)
By substituting (39) into (38), we obtain an upper bound of the Rademacher complexity $\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$, as follows:
@@ B-3 PAC upper bound of the excess cost
Lemma 2 (Generalization bound via uniform convergence [42])
@@
B-4 Extension to other types of NNs
The performance guarantee established for L2-regularized NNs can be extended to other types of NNs with minor modifications to the derivation process. Taking L1-regularized NNs as an example, the L1-regularization constraint implies:
$\|\mathbf{w}^{k}_{j}\|_{1}\leq W^{\mathrm{max}}\quad\Rightarrow\quad\|\mathbf{w}^{k}_{j}\|_{\infty}\leq\|\mathbf{w}^{k}_{j}\|_{1}\leq W^{\mathrm{max}}.$ (42)
Following Eqs. (32)-(34), the empirical Rademacher complexity ^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$ is bounded by:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$
$\leq\frac{(W^{\mathrm{max}})^{2}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \|\overline{\mathbf{w}}\|_{2}=1\end{subarray}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|\right].$ (43)
Further applying Eqs. (35)–(38), we derive:
$\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})\leq\sqrt{2}C_{\ell}\cdot(2W^{\mathrm{max}})^{K}\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1,L1}),$ (44)
where the hypothesis class is defined as ¯$\overline{\mathcal{H}}^{1,L1}=\{\mathbf{x}\mapsto\overline{\mathbf{w}}^{\intercal}\mathbf{x}:\|\overline{\mathbf{w}}\|_{1}\leq 1\}$. According to [41, Theorem 3], its Rademacher complexity is bounded by:
$\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1,L1})\leq X^{\mathrm{max}}\sqrt{\frac{\ln p}{M}},$ (45)
where $p$ is the dimensionality of the observable feature vector $\mathbf{x}$. The remaining steps in the derivation directly follow the process used for L2-regularized NNs to establish the performance guarantee for L1-regularized NNs.
References [43, 44, 45] provide bounds for the empirical Rademacher complexity ^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$ when RNNs or CNNs are used as the hypothesis classes for prediction. Substituting these bounds into (29) yields the Rademacher complexity bound for the loss class $\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$. By further substituting this bound into (41), we can obtain the extension of the performance guarantee to RNNs and CNNs.
Appendix C
C-1 Extension to the AC formulation
After replacing the DC power flow model in (2b)-(2c) with the AC formulation, the second-stage problem becomes non-convex. Nevertheless, convex relaxations of the AC power flow model have been proposed [46, 47]. According to the envelope theorem mentioned in Proposition 1, the gradient of an optimization problem’s Lagrangian at saddle points with respect to certain parameters can be explicitly expressed using dual variables. If our second-stage problem has a tight convex relaxation, its Lagrangian at the saddle points will equal its optimal objective value. Therefore, the gradient of the RLD loss can be computed using dual variables, enabling effective NN training. With this tight convex relaxation, the value function $Q(\mathbf{u},\mathbf{d})$ remains Lipschitz continuous, as shown in [37, Theorem 5.2]. Thus, the RLD loss also retains its Lipschitz continuity. By replacing the constant $C_{\ell}$ in (12) with the Lipschitz constant of the new $Q(\mathbf{u},\mathbf{d})$, we can apply the same procedure detailed in Appendix B to establish the generalization performance guarantee for the AC case.
C-2 Extension to incorporating multiple periods and additional operation constraints
When ramping constraints, production limits, and generation reduction costs are considered, the RLD problem becomes
$\min_{\mathbf{u}_{t},\forall t\in\mathcal{T}}\quad\left(\sum_{t\in\mathcal{T}}\bm{\alpha}_{t}^{\intercal}\mathbf{u}_{t}\right)+\mathbb{E}_{\mathbf{d}\sim\mathcal{D}(\mathbf{x})}[Q(\mathbf{u},\mathbf{d})],$ (46)
where we introduce a temporal dimension with time intervals indexed by $t\in\mathcal{T}$ to account for ramping constraints. The second-stage cost $Q(\mathbf{u},\mathbf{d})$ represents the real-time dispatch cost given the trajectories of first-stage decision $\mathbf{u}=[\mathbf{u}_{t},\forall t\in\mathcal{T}]$ and realized net demand $\mathbf{d}=[\mathbf{d}_{t},\forall t\in\mathcal{T}]$:
$Q(\mathbf{u},\mathbf{d}$ ))=) =
$\min_{(\mathbf{g}_{t},\bm{\theta}_{t})_{t\in\mathcal{T}}}$ $\sum_{t\in\mathcal{T}}\left((\bm{\beta}_{t}^{+})^{\intercal}\mathbf{g}_{t}^{+}+(\bm{\beta}_{t}^{-})^{\intercal}\mathbf{g}_{t}^{-}\right),$ (47a)
s.t. $\mathbf{g}_{t}^{+}-\mathbf{g}_{t}^{-}=\mathbf{g}_{t},\mathbf{g}_{t}^{+}\geq\bm{0},\mathbf{g}_{t}^{-}\geq\bm{0},\ \forall t\in\mathcal{T},$ (47b)
$\underline{\mathbf{u}}\leq\mathbf{u}_{t}+\mathbf{g}_{t}\leq\overline{\mathbf{u}},\ \forall t\in\mathcal{T},$ (47c)
$-\mathbf{r}^{\mathrm{down}}\leq(\mathbf{u}_{t}+\mathbf{g}_{t})-(\mathbf{u}_{t-1}+\mathbf{g}_{t-1})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\leq\mathbf{r}^{\mathrm{up}},\ \forall t\in\mathcal{T},$ (47d)
$\mathbf{u}_{t}+\mathbf{g}_{t}-\mathbf{d}_{t}=\mathbf{B}\bm{\theta}_{t},\ \forall t\in\mathcal{T},$ (47e)
$-\mathbf{f}^{\mathrm{max}}\leq\mathbf{F}\bm{\theta}_{t}\leq\mathbf{f}^{\mathrm{max}},\ \forall t\in\mathcal{T},$ (47f)
where $\mathbf{g}_{t}^{+}$ and $\mathbf{g}_{t}^{-}$ represent real-time generation adjustments, with costs captured by $\bm{\beta}_{t}^{+}$ and $\bm{\beta}_{t}^{-}$ for each time period $t$. Constraints (47c) and (47d) enforce production and ramping limits, while (47e) and (47f) describe power flow and branch capacity limits. All operational constraints are imposed in the second stage, ensuring that real-time generation adjustments are feasible and implementable.
Since the second-stage problem remains linear, the gradient $\nabla_{\bm{\phi}}Q_{t}$ can be explicitly expressed using dual variables according to Proposition 1, allowing for effective training. The generalization performance guarantee can also be derived in the same manner as outlined in Appendix B, as the linear structure of the second-stage problem remains unchanged.